OpenStudy (anonymous):

The population of a town decreased from 217,000 to 213,000 in 2 years. Assuming this trend continues what will the pop be in 3 additional years??

OpenStudy (anonymous):

12.000

OpenStudy (campbell_st):

is a linear function or exponential growth and decay..

OpenStudy (anonymous):

Difference of 217,000 and 213,000 is 4,000. So the population decreased by 4000 in two years to 213,000 people. So for three additional years it will decrease by 213,000 – (4000+2000) = 207,000

OpenStudy (campbell_st):

I'll assume exponential find the decay constant \[P _{0} = 217000, t = 2\] \[ A = P _{0}e^{-kt}\]\ \[213000 = 217000e^{-2k}\] \[-2k = \ln(213000/217000)\] k = 0.009303 so the equation is \[A = 217000e^{-0.009303t}\] substitute t = 5 and solve to find the population